Skip to content

OSG Technology Area Meeting, 14 May 2018

Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Carl, Derek, Marian, Mat, TimT

Announcements

  • HEPiX this week: Edgar will have limited availability
  • Suchandra driving down to IU to pick up servers today
  • HTCondor week next week: Madison team and Edgar will have limited availability

Triage Duty

  • This week: Carl
  • Next week: Edgar
  • 13 (+2) open tickets
  • 0 (+0) open host certificate tickets
  • 11 (+11) open user certificate tickets
  • 8 (-2) open LCMAPS VOMS tickets

JIRA

# of tickets Δ State
109 -1 Open
40 +0 In Progress
7 -17 Ready for Testing
1 -10 Ready for Release

OSG Software Team

  • JIRA migration: This Wednesday 5/16 9am CDT
  • Host certificate document is broken
    • AI (Carl): Review formatting issues and duplicate sub-sections
  • Derek fixed a bad reference to myosg.opensciencegrid.org in osg-display-data
    • AI (Carl): Cut a new build for the GOC repo
  • Triage Duty
    • AI (TimT): revisit unattended tickets from last week
    • Derek: What were the results of the Freshdesk/HCC RT test?
    • Remember to monitor user and host certificate tickets in Footprints
    • AI (everyone): Update all of your Freshdesk tickets appropriately,
      either respond or set the status to Waiting on Customer Response.
    • AI (everyone): Let's give LCMAPS VOMS tickets one last ping and if we don't get a response, close them next week.
  • Office hours: Tuesdays 2-3pm CDT and Thursdays 10-11am CDT

Discussion

  • IU repo host retirement was delayed to May 23
  • JIRA DNS will be updated but if there are delays in propagation, you can access JIRA via opensciencegrid.atlassian.net
  • A location for JIRA backups need to be found. AFS?
  • AI (BrianL): Put up note about canceling next week's office hours
  • AI (Derek): Test Freshdesk + HCC RT this week

Support Update

UFL (Derek) - Assisted with April accounting records issuess due to multiple CEs

OSG Release Team

3.4.13 Δ Status
11 +11 Open
9 +9 In Progress
4 +4 Ready for Testing
0 +0 Ready for Release
24 +24 Total
  • GOC
    • Testing
      • mash-scripts: grid.iu.edu becomes opensciencegrid.org
      • osg-display: grid.iu.edu becomes opensciencegrid.org
      • repo-update-cadist: RPM packaging
    • Ready for Release
      • Package repo.opensciencegrid.org scripts
      • rev-perfsonar: grid.iu.edu becomes opensciencegrid.org
  • Data
    • IGTF 1.91
  • OSG 3.4.13
    • Testing
      • voms-proxy-direct
      • Pegasus 4.8.2
      • Singularity 2.5.1
      • OWAMP 3.5.6
    • Ready for Release
      • Nothing Yet

Discussions

None this week

OSG Investigations Team

GOC Transition is going to dominate the Investigations Team's time over the next couple of weeks. I want to give general status of the transitions we are participating in. So, I will give a couple of states:

  1. Have not started
  2. In development - Could be provisioning hardware, adapting software, installing software…
  3. Ready for testing - In a state where it can be tested by the consumers
  4. Deployed
Service Status
OIM In Development
OASIS Deployed
CVMFS Deployed
Software Repo Deployed
OSG Display Deployed
StashCache Redirector Deployed
OSG Collector Ready for testing

Changes:

  • Will start social compaign to have people restart their CEs to pick up the DNS change for the new collector.

Ongoing

Discussions

None this week